OpenStudy (melodyeason74):

complete the square / find the missing term z^2 - 2z

OpenStudy (jadeishere):

In order to use the "Complete the Square" it needs to be a perfect square trinomial, I believe....

OpenStudy (melodyeason74):

the question says to simplify and write as integer or a fraction

OpenStudy (jadeishere):

Um, I'm confused, and I'm not sure I can help. I need to be able to see the entire question.. Sorry, maybe @alivejeremy can help:)

OpenStudy (melodyeason74):

ah, that was the whole question... thanks anyway though

OpenStudy (alivejeremy):

There are no terms

OpenStudy (melodyeason74):

\[z ^{2}-2z\]

OpenStudy (alivejeremy):

z(z−2)

OpenStudy (melodyeason74):

\[z ^{2}-2z+?\] this is what the question shows, all it says is find the missing term that complete the square, ​(Simplify your answer. Type an integer or a​ fraction.)

OpenStudy (alivejeremy):

=z^2−2z

OpenStudy (alivejeremy):

There shouldn't be nothing missing.

OpenStudy (alivejeremy):

Wait let me check again.

OpenStudy (austin.l):

\[z^2+2z+x=0\] That seems like how I would start it, solve for it from there. But I'm a bit rusty and I may be wrong.

OpenStudy (austin.l):

Correction, \[z^2+2z-x=0\]

OpenStudy (melodyeason74):

ok, i will try that

OpenStudy (alivejeremy):

Do you want us to find the degree ?

OpenStudy (alivejeremy):

Because you really can't simplify

OpenStudy (alivejeremy):

and there's really nothing missing.

OpenStudy (alivejeremy):

Unless you want us the find what make =z^2−2z ?

OpenStudy (melodyeason74):

um... sure if thats how its supposed to be solved, i wrote exactly what was on my paper so i dont know

OpenStudy (austin.l):

I mean, to complete the square it is usually an expression set equal to 0. I'm not entirely sure how to tackle this to be honest. Are you sure that is the entire problem?

OpenStudy (alivejeremy):

(z-1)^2 - 1

OpenStudy (melodyeason74):

im positive, that is why im so confused

OpenStudy (welshfella):

you can wrte it in the form (z - 1)^2 - 1

OpenStudy (alivejeremy):

What I was thinking ^

OpenStudy (welshfella):

the question is confusing

OpenStudy (welshfella):

maybe the missing term is -1 ???

OpenStudy (abdullahm):

a^2 + 2ab + b^2 = (a + b)^2 We want to find the term "b^2" and so we can reduce it into the form of (a + b)^2 The way to find b^2 is by taking 2ab dividing it by 2a and squaring it. In our case we need to do \(\Large \left(\frac{2z}{2z}\right)^2 = 1\) And that is our missing term.

OpenStudy (austin.l):

But that still doesn't alleviate the issue of us having to complete the square when all we have is an expression and not an equation. Because if that expression was set equal to zero, I think I might do something along the lines of: \[z(2+z)=0\]

OpenStudy (abdullahm):

Completing the square is a way of adding a term to be able to factor it into a perfect square trinomial.

OpenStudy (welshfella):

yea I think abdullah is right the missing term is + 1 that makes z^2 - 2z + 1 a perfect square

OpenStudy (melodyeason74):

thanks y'all, sorry for the trouble

OpenStudy (welshfella):

there is no need to equate it to zero

OpenStudy (austin.l):

Oh my goodness, I was using a + instead of a - the entire time. My way would have worked if I had had the signs correct.

OpenStudy (welshfella):

yw

OpenStudy (austin.l):

Long way around it, but good job @AbdullahM Take your filthy medal sir.

OpenStudy (abdullahm):

The last term in a perfect square trinomial will always be positive. (a + b)^2 = a^2 + 2ab + b^2 (a - b)^2 = a^2 - 2ab + b^2